OK THIS IS A DUMB QUESTION BUT I HAVE NO ONE TO ANWSER IT RIGHT NOW. WHEN I'M ASKED TO FIND THE DISTANCE OF TWO POINTS THAT MEANS I FIND THE SLOPE RIGHT?
Answer by venugopalramana(3286)

OK THIS IS A DUMB QUESTION BUT I HAVE NO ONE TO ANWSER IT RIGHT NOW. WHEN I'M ASKED TO FIND THE DISTANCE OF TWO POINTS THAT MEANS I FIND THE SLOPE RIGHT?
NO IF THE POINTS ARE A WITH (X1,Y1)...SAY (1,2) AND B WITH (X2,Y2)SAY (4,6) AS COORDINATES,THEN
AB=DISTANCE BETWEEN 2 POINTS THAT IS FROM A TO B .IT IS GIVEN BY
SQRT[(X2-X1)^2 +(Y2-Y1)^2]=SQRT[(4-1)^2+(6-2)^2]=SQRT(9+16)=5
SLOPE OF AB IS THE TANGENT OF THE ANGLE MADE BY AB WITH X AXIS.IT IS GIVEN BY
(Y2-Y1)/(X2-X1)=(6-2)/(4-1)=4/3=TAN(THETA),WHERE THETA IS ANGLE BETWEEN AB AND X AXIS